OpenStudy (anonymous):

Factor completely: 5ab + 3ay + 5b + 3y

OpenStudy (anonymous):

a( 5b + 3y) _ 5b +4y = (5b + 3y) (a-1)

OpenStudy (anonymous):

5b (a +1)+ 3y (a+1)= (a+1)(5b+3y) another approach